Suppose that you have 72 F2 flies from a cross, and expected that they would show a phenotype ratio of: 9 red eye/normal wing please please answer this
Answer #1

According to the question, 9 out of 16 offsprings are with red eyes and normal wings. So, out of 72 F2 flies, their number would be = (9/16)×72= 0.5625×72 = 40.5.

Answer. 40.5

Thanks